# Negative Dependence as a toolbox for machine learning : review and new developments

### v1

@@ 6.4.2 PAC bound
$\frac{1}{N}|L(f)|\geq c,\quad\text{ for some }c>0\text{ uniformly over }\mathcal{F}.$ (A.0)
@@
=$\mathcal{F}=\{f_{\theta}:\theta\in\Theta\},\text{ where $\Theta$ is a bounded subset of $\mathbb{R}^{D}$ for some $D$}.$ (A.2)
@@
In these two common situations, we can perform a chaining argument to obtain the following result (see [BGST24]).
@@
$\mathbb{P}\Big{(}\exists f\in\mathcal{F}:\Big{|}\frac{L_{\mathcal{S}}(f)}{L(f)}-1\Big{|}\geq\varepsilon\Big{)}\leq 2\exp\Big(6D-{c^{2}\varepsilon^{2}\over 16AV}\Big{missing}),\quad 0\leq\varepsilon\leq{4A\rho mV\over 3c\sup_{f\in\mathcal{F}}\|f\|_{\infty}}\cdot$
@@
[…]
@@
For independent sampling, ${\mathop{\mathrm{Var}}\nolimits}[N^{-1}L_{\mathcal{S}}(f)]$ is of order $O(m^{-1})$. In contrast, sampling with DPPs often results in a variance of order $O(m^{-(1+\delta)})$ for some $\delta>0$ (e.g., see [BH20, BGL21]). Therefore, the range of $\varepsilon$ for which we can apply our concentration result is $O(m^{-\delta})$. Substituting $\varepsilon=m^{-\alpha}(\alpha\geq\delta)$ into Theorem 6.13 yields the upper bounds $2\exp(6D-C^{\prime}m^{1+\delta-2\alpha})$ and $2\exp(CD+\alpha D\log m-C^{\prime}m^{1+\delta-2\alpha})$, where $C$ and $C^{\prime}$ are positive constants independent of $m$ and $N$. Both of them converge to 00 as $m\rightarrow\infty$, provided $\alpha<(1+\delta)/2$. Hence, the accuracy rate $\varepsilon$ can be as small as $m^{-1/2-\delta^{\prime}/2}$, for any $0<\delta^{\prime}<\delta$, which clearly outperforms the best accuracy rate of $m^{-1/2}$ for i.i.d. sampling.
The assumptions in Theorem 6.13 are fairly moderate and reasonable. For i.i.d. sample $\mathcal{S}$ with expected size $m$, the probability $\mathbb{P}(x\in\mathcal{S})=m/N$ applies to each $x\in\mathcal{X}$. For a DPP sample $\mathcal{S}$ with kernel $\mathbf{K}$, we have $\mathbb{P}(x_{i}\in\mathcal{S})=\mathbf{K}_{ii}$. Therefore, assuming $\mathbf{K}_{ii}\geq\rho\cdot m/N$ for each $i$ and some $\rho>0$ ensures that every point in the dataset $\mathcal{X}$ has a fair chance of being selected. This also helps prevent the estimated loss $L_{\mathcal{S}}(f)=\sum_{x_{i}\in\mathcal{S}}f(x_{i})/\mathbf{K}_{ii}$ from blowing up to infinity. In the second scenario, the assumption $|\mathcal{S}|\leq B\cdot m$ a.s. for some $B>0$ is not strictly required but is included to simplify the presentation of the results. Indeed, for a DPP $\mathcal{S}$ with Hermitian kernel, it is known that $|\mathcal{S}|$ is equal in distribution to the sum of independent Bernoulli variables, which consequently implies that $|\mathcal{S}|$ highly concentrates around its mean $m$. This property allows us to relax the assumption on the size of $|\mathcal{S}|$ by applying a conditioning argument. However, the assumption $|\mathcal{S}|\leq B\cdot m$ a.s. holds for most kernels of interest; DPPs with projection kernels being typical and significant examples. In machine learning terms, this means that the coresets are not much bigger than their expected size $m$; whereas in practice, sampling methods often produce coresets of a fixed size, as observed with projection DPPs. We refer the readers to [BGST24] for more detailed discussion.
@@ 7.1 Background
This area is also intertwined closely with the stochastic geometry of point processes. The stochastic geometry of negatively dependent point processes has attracted considerable attention in recent years in the applied probability community, including but not limited to the problems of continuum percolation or the Gilbert disk model [RMIY07, BB10, BHKM18, MS14, GKP16].
@@ 7.2 Stochastic Networks and hyperuniformity
The classical and most popularly studied model for stochastic networks is based on the Poisson point process [BB10, BHKM18, Hae12]. However, Poissonian networks exhibit similar issues to independent random points, thereby limiting their performance.
In view of this, dependent point fields have been investigated in this area; most notably the so-called Ginibre network which is based on the Ginibre ensemble – a DPP that is intimately related to random matrices and 2D Coulomb gases. In the celebrated work [MS14], the authors proposed a stochastic network based on the eigenvalues from the Ginibre ensemble, and demonstrated rigorously that such a network provides provably improved performance over standard Poissonian networks. To this end, they observed that for quantities that depend solely on the absolute values of the Ginibre points, such as the coverage probability or link success probability, we can take advantage of the fact that the process of absolute values can be equivalently described by independent gamma random variables.
@@
Hyperuniformity refers to the suppression of fluctuations in particle numbers in a spatial point process. In Poissonian systems, the variance in the number of points within a large spatial domain increases proportionally to the volume (a phenomenon described in physics as extensive fluctuations). However, in hyperuniform systems, these fluctuations are of a smaller order relative to the volume (sub-extensive), often scaling only with the surface area of the domain. Such behaviour manifests itself in a wide array of natural systems [HKPV09, GL16a, Tor18, TZS15, GL18], and allows for elegant and effective interpretations in terms of their power spectra [TS03, BKM14, AGL20]. We refer the readers to [GL16a, Cos21] for a detailed discussion on hyperuniformity, and to [HGBLR23] for statistical tests of hyperuniformity for point processes.
@@ 7.3 Stochastic networks based on disordered lattices
In [GMS22], the authors investigate disordered lattice networks, aiming to find a best-of-both-worlds compromise between Poissonian and Ginibre networks. As the name indicates, these networks are based on point processes that are i.i.d. random perturbations of Euclidean lattices. It is known that under very general conditions – as soon as the tail of the perturbing random variable decays faster that $\|x\|^{-d}$ for a $d$-dimensional lattice – the resulting disordered lattice constitutes a hyperuniform process [GL16a, GL16b].
@@
where $X_{B}$ is the location of the node of the network nearest to the origin, $d$ is the ambient dimension, and $\beta>0$ is a fixed modelling parameter related to the decay of the so-called path loss function [BB10, MS14].
Focusing on Gaussian disorder for the perturbed lattice networks, the authors in [GMS22] demonstrate that for an appropriate choice of the disorder strength, the coverage probabilities roughly match those of a Ginibre random network, thereby achieving the best of both worlds phenomenon aspired for.
Interestingly, at the same disorder strength, the perturbed lattice point process appears to be the closest to the Ginibre random point field. In order to achieve this comparison, [GMS22] introduces an approach towards robust measurement of the “distance” between point processes by a quantitative comparison of their persistence homology, a technique that we expect to see increasing usefulness in coming years.
In addition to elaborate theoretical analysis demonstrating power law asymptotics for the coverage probability of disordered lattice networks and of interpolation properties discussed above, [GMS22] obtains stochastic approximations to the SINR variable in various limiting regimes of interest. An interesting finding in this vein is the emergence of the Epstein Zeta Function of the underlying lattice as a key determinant of network statistics. This leads to the optimal choices of lattices as the minimizers of this lattice energy function, which is given by the triangular lattice in 2D and face centered cubic (FCC) lattice in 3D.
@@ 8 Spectrogram analysis and Gaussian Analytic Functions
In time-frequency analysis, Short Time Fourier Transforms (abbrv. STFTs) are foundational objects [Gr 01, Fla98, Coh95]. Given a signal $f\in L^{2}(\mathbb{R})$ and a window function $\phi\in L^{2}(\mathbb{R})$, the STFT $\mathcal{F}_{\phi}f(\cdot,\cdot)$ of $f$ with respect to $\phi$ is defined as the inner product between f and shifted $\phi$ (in the sense of time and frequency). More precisely,
@@
We will particularly choose $\phi(x)=2^{1/4}e^{-\pi x^{2}}$ as in [BFC20, GLS22], to make use of the Bargmann transform, which will be introduced later. In this setting, the STFT is also known as the Gabor transform [Gr 01].
@@
Spectrogram analysis is highly effective in many applications [Gr 01, Fla15]. Classically, the maxima of the spectrogram have been extensively studied. This is related to the understanding that these capture greater energy from the spectrogram and, therefore, provide more information about the signal [Gr 01, Fla98, Coh95, Fla18]. However, in recent years, a different approach to spectrogram analysis has been investigated, focusing on the zeros of the spectrogram instead of the maxima. This new line of investigation originated from a seminal paper by Flandrin [Fla15].
In relation to this direction, Bardenet et al. [BFC20] observed that the zero set of the spectrogram of the Gaussian white noise has the same statistical distribution as the zero set of the planar Gaussian Analytic Function (GAF). In [BH21], Bardenet and Hardy further explored the connection between several time-frequency transforms of Gaussian white noises and different models of GAFs (planar, elliptic, hyperbolic). In particular, the STFT with respect to a Gaussian window corresponds to the planar GAF, and the analytic wavelet transform [DP88] is related to the hyperbolic GAF. Thus, spectrogram analysis is linked to the theory of Gaussian Analytic Functions, a rich and elegant topic of independent mathematical interest [HKPV09, PV05, Sod00].
@@ 8.1 Zeros of Gaussian Analytic Functions
Definition 8.2 (Gaussian analytic functions, [HKPV09]).
@@
Figure 3: Zeros of 3 canonical models of GAFs (cf. [BH21])
These 3 models are characterized by the property that the zero sets are invariant (in distribution) and ergodic under the action of the group of isomorphisms of the ambient domains (see [HKPV09], Chapter 2).
The behavior of the zeros of GAFs is quite similar to that of determinantal point processes (DPPs), as both exhibit the so-called hyperuniformity in the spatial distribution of points, which is of independent interest in statistical physics [GL16b, GL16a, TS03]. In fact, for the hyperbolic GAF with $L=1$, it has been shown by Peres and Virag [PV05] that the zeros of $f_{1}$ is a DPP on $\mathbb{D}$ with the Bergman kernel $K(z,w)=\pi^{-1}(1-z\bar{w})^{-2}$ and the background measure is the Lebesgue measure on $\mathbb{D}$. The zeros of GAFs also exhibit repulsiveness, but at short ranges. In particular, the probability of having 2 zeros in a same small disk of radius $\varepsilon$ is of order $\mathcal{O}(\varepsilon^{6})$ (see [HKPV09], Chapter 3). In contrast, from (2.2) with $d=2$, that probability for Poisson point processes is of order $\varepsilon^{4}$. For more details about GAFs, we refer the readers to [HKPV09, Sod00].
@@ 8.2.1 Bargmann transform and Hermite functions
$\mathcal{B}f(z)=2^{1/4}\int_{\mathbb{R}}f(x)\exp\Big(2\pi xz-\pi x^{2}-\frac{\pi}{2}z^{2}\Big{missing})\mathrm{d}x,\quad z\in\mathbb{C}.$
@@
$\mathcal{F}_{\phi}f(z)=\exp\Big(-\pi iuv-\frac{\pi}{2}|z|^{2}\Big{missing})\mathcal{B}f(\bar{z}),$
@@ 8.3 The short-time Fourier transforms of white noise
Using Bargmann transform, the STFT of the white noise $\xi$ is then given by the following theorem (see [BFC20, GLS22]).
@@
$\mathcal{F}_{\phi}\xi(z)=\sqrt{\pi}\exp\Big(i\pi uv-\frac{\pi}{2}|z|^{2}\Big{missing})\sum_{k=0}^{\infty}\langle\xi,h_{k}\rangle\sqrt{\frac{\pi^{k}}{k!}}z^{k},$ (8.3)
@@
Since $\langle\xi,h_{k}\rangle$ are i.i.d. real Gaussians, this is also called the symmetric planar Gaussian analytic function. However, the zero set in this case is not stationary, which poses an obstacle in using techniques from spatial statistics. Motivated by early results by Prosen, the authors of [BFC20] argue that the zero set of the spectrogram of real Gaussian white noise could be effectively approximated by the zero set of the spectrogram of complex Gaussian white noise. It turns out that the latter is precisely the zero set of the planar GAF, which has been shown to be invariant under isomorphisms of the complex plane.

### v4

@@ 6.4.2 PAC bound
[…]
@@
[…]
@@
In these two common situations, we can perform a chaining argument to obtain the following result (see [NIPS2024]).
@@
$\mathbb{P}\Big(\exists f\in\mathcal{F}:\Big|\frac{L_{\mathcal{S}}(f)}{L(f)}-1\Big|\geq\varepsilon\Big)\leq 2\exp\Big(6D-{c^{2}\varepsilon^{2}\over 16AV}\Big.),\quad 0\leq\varepsilon\leq{4A\rho mV\over 3c\sup_{f\in\mathcal{F}}\|f\|_{\infty}}\cdot$
@@
$\mathbb{P}\Big(\exists f\in\mathcal{F}:\Big|\frac{L_{\mathcal{S}}(f)}{L(f)}-1\Big|\geq\varepsilon\Big)\leq 2\exp\Big(CD-D\log\varepsilon-{c^{2}\varepsilon^{2}\over 16AV}\Big.),\>0\leq\varepsilon\leq{4A\rho mV\over 3c\sup_{f\in\mathcal{F}}\|f\|_{\infty}}\cdot$
@@
For independent sampling, ${\mathop{\mathrm{Var}}\nolimits}[N^{-1}L_{\mathcal{S}}(f)]$ is of order $O(m^{-1})$. In contrast, sampling with DPPs often results in a variance of order $O(m^{-(1+\delta)})$ for some $\delta>0$ (e.g., see [OPE-AOAP, OPE-NIPS]). Therefore, the range of $\varepsilon$ for which we can apply our concentration result is $O(m^{-\delta})$. Substituting $\varepsilon=m^{-\alpha}(\alpha\geq\delta)$ into Theorem 6.13 yields the upper bounds $2\exp(6D-C^{\prime}m^{1+\delta-2\alpha})$ and $2\exp(CD+\alpha D\log m-C^{\prime}m^{1+\delta-2\alpha})$, where $C$ and $C^{\prime}$ are positive constants independent of $m$ and $N$. Both of them converge to 0 as $m\rightarrow\infty$, provided $\alpha<(1+\delta)/2$. Hence, the accuracy rate $\varepsilon$ can be as small as $m^{-1/2-\delta^{\prime}/2}$, for any $0<\delta^{\prime}<\delta$, which clearly outperforms the best accuracy rate of $m^{-1/2}$ for i.i.d. sampling.
The assumptions in Theorem 6.13 are fairly moderate and reasonable. For i.i.d. sample $\mathcal{S}$ with expected size $m$, the probability $\mathbb{P}(x\in\mathcal{S})=m/N$ applies to each $x\in\mathcal{X}$. For a DPP sample $\mathcal{S}$ with kernel $\mathbf{K}$, we have $\mathbb{P}(x_{i}\in\mathcal{S})=\mathbf{K}_{ii}$. Therefore, assuming $\mathbf{K}_{ii}\geq\rho\cdot m/N$ for each $i$ and some $\rho>0$ ensures that every point in the dataset $\mathcal{X}$ has a fair chance of being selected. This also helps prevent the estimated loss $L_{\mathcal{S}}(f)=\sum_{x_{i}\in\mathcal{S}}f(x_{i})/\mathbf{K}_{ii}$ from blowing up to infinity. In the second scenario, the assumption $|\mathcal{S}|\leq B\cdot m$ a.s. for some $B>0$ is not strictly required but is included to simplify the presentation of the results. Indeed, for a DPP $\mathcal{S}$ with Hermitian kernel, it is known that $|\mathcal{S}|$ is equal in distribution to the sum of independent Bernoulli variables, which consequently implies that $|\mathcal{S}|$ highly concentrates around its mean $m$. This property allows us to relax the assumption on the size of $|\mathcal{S}|$ by applying a conditioning argument. However, the assumption $|\mathcal{S}|\leq B\cdot m$ a.s. holds for most kernels of interest; DPPs with projection kernels being typical and significant examples. In machine learning terms, this means that the coresets are not much bigger than their expected size $m$; whereas in practice, sampling methods often produce coresets of a fixed size, as observed with projection DPPs. We refer the readers to [NIPS2024] for more detailed discussion.
@@ 7.1 Background
This area is also intertwined closely with the stochastic geometry of point processes. The stochastic geometry of negatively dependent point processes has attracted considerable attention in recent years in the applied probability community, including but not limited to the problems of continuum percolation or the Gilbert disk model [yogesh, Baccelli-Blasz, Blaszczyszyn_Haenggi_Keeler_Mukherjee_2018, shirai-miyoshi, ghosh_aop].
@@ 7.2 Stochastic Networks and hyperuniformity
The classical and most popularly studied model for stochastic networks is based on the Poisson point process [Baccelli-Blasz, Blaszczyszyn_Haenggi_Keeler_Mukherjee_2018, Haenggi]. However, Poissonian networks exhibit similar issues to independent random points, thereby limiting their performance.
In view of this, dependent point fields have been investigated in this area; most notably the so-called Ginibre network which is based on the Ginibre ensemble – a DPP that is intimately related to random matrices and 2D Coulomb gases. In the celebrated work [shirai-miyoshi], the authors proposed a stochastic network based on the eigenvalues from the Ginibre ensemble, and demonstrated rigorously that such a network provides provably improved performance over standard Poissonian networks. To this end, they observed that for quantities that depend solely on the absolute values of the Ginibre points, such as the coverage probability or link success probability, we can take advantage of the fact that the process of absolute values can be equivalently described by independent gamma random variables.
@@
Hyperuniformity refers to the suppression of fluctuations in particle numbers in a spatial point process. In Poissonian systems, the variance in the number of points within a large spatial domain increases proportionally to the volume (a phenomenon described in physics as extensive fluctuations). However, in hyperuniform systems, these fluctuations are of a smaller order relative to the volume (sub-extensive), often scaling only with the surface area of the domain. Such behaviour manifests itself in a wide array of natural systems [hough2009zeros, GhoshLebowitz, TORQUATO, torquato2, ghosh-lebowitz-cmp], and allows for elegant and effective interpretations in terms of their power spectra [PhysRevE, baake, ghoshlebowitz-cmp2]. We refer the readers to [GhoshLebowitz, coste] for a detailed discussion on hyperuniformity, and to [lacieze_hyperuniform] for statistical tests of hyperuniformity for point processes.
@@ 7.3 Stochastic networks based on disordered lattices
In [ghosh-shirai], the authors investigate disordered lattice networks, aiming to find a best-of-both-worlds compromise between Poissonian and Ginibre networks. As the name indicates, these networks are based on point processes that are i.i.d. random perturbations of Euclidean lattices. It is known that under very general conditions – as soon as the tail of the perturbing random variable decays faster that $\|x\|^{-d}$ for a $d$-dimensional lattice – the resulting disordered lattice constitutes a hyperuniform process [GhoshLebowitz, Ghosh2016NumberRI].
@@
where $X_{B}$ is the location of the node of the network nearest to the origin, $d$ is the ambient dimension, and $\beta>0$ is a fixed modelling parameter related to the decay of the so-called path loss function [Baccelli-Blasz, shirai-miyoshi].
Focusing on Gaussian disorder for the perturbed lattice networks, the authors in [ghosh-shirai] demonstrate that for an appropriate choice of the disorder strength, the coverage probabilities roughly match those of a Ginibre random network, thereby achieving the best of both worlds phenomenon aspired for.
Interestingly, at the same disorder strength, the perturbed lattice point process appears to be the closest to the Ginibre random point field. In order to achieve this comparison, [ghosh-shirai] introduces an approach towards robust measurement of the “distance” between point processes by a quantitative comparison of their persistence homology, a technique that we expect to see increasing usefulness in coming years.
In addition to elaborate theoretical analysis demonstrating power law asymptotics for the coverage probability of disordered lattice networks and of interpolation properties discussed above, [ghosh-shirai] obtains stochastic approximations to the SINR variable in various limiting regimes of interest. An interesting finding in this vein is the emergence of the Epstein Zeta Function of the underlying lattice as a key determinant of network statistics. This leads to the optimal choices of lattices as the minimizers of this lattice energy function, which is given by the triangular lattice in 2D and face centered cubic (FCC) lattice in 3D.
@@ 8 Spectrogram analysis and Gaussian Analytic Functions
In time-frequency analysis, Short Time Fourier Transforms (abbrv. STFTs) are foundational objects [TFA_Grochenig, TFA_Flandrin, TFA_Cohen]. Given a signal $f\in L^{2}(\mathbb{R})$ and a window function $\phi\in L^{2}(\mathbb{R})$, the STFT $\mathcal{F}_{\phi}f(\cdot,\cdot)$ of $f$ with respect to $\phi$ is defined as the inner product between f and shifted $\phi$ (in the sense of time and frequency). More precisely,
@@
We will particularly choose $\phi(x)=2^{1/4}e^{-\pi x^{2}}$ as in [BARDENET_Spec, GhoshGAF], to make use of the Bargmann transform, which will be introduced later. In this setting, the STFT is also known as the Gabor transform [TFA_Grochenig].
@@
Spectrogram analysis is highly effective in many applications [TFA_Grochenig, Spectrogram_Flandrin]. Classically, the maxima of the spectrogram have been extensively studied. This is related to the understanding that these capture greater energy from the spectrogram and, therefore, provide more information about the signal [TFA_Grochenig, TFA_Flandrin, TFA_Cohen, Flandrin_2018]. However, in recent years, a different approach to spectrogram analysis has been investigated, focusing on the zeros of the spectrogram instead of the maxima. This new line of investigation originated from a seminal paper by Flandrin [Spectrogram_Flandrin].
In relation to this direction, Bardenet et al. [BARDENET_Spec] observed that the zero set of the spectrogram of the Gaussian white noise has the same statistical distribution as the zero set of the planar Gaussian Analytic Function (GAF). In [BARDENET_GAF], Bardenet and Hardy further explored the connection between several time-frequency transforms of Gaussian white noises and different models of GAFs (planar, elliptic, hyperbolic). In particular, the STFT with respect to a Gaussian window corresponds to the planar GAF, and the analytic wavelet transform [Daubechies_1988] is related to the hyperbolic GAF. Thus, spectrogram analysis is linked to the theory of Gaussian Analytic Functions, a rich and elegant topic of independent mathematical interest [hough2009zeros, PeresVirag, sodin].
@@ 8.1 Zeros of Gaussian Analytic Functions
Definition 8.2 (Gaussian analytic functions, [hough2009zeros]).
@@
Figure 3: Zeros of 3 canonical models of GAFs (cf. [BARDENET_GAF])
These 3 models are characterized by the property that the zero sets are invariant (in distribution) and ergodic under the action of the group of isomorphisms of the ambient domains (see [hough2009zeros], Chapter 2).
The behavior of the zeros of GAFs is quite similar to that of determinantal point processes (DPPs), as both exhibit the so-called hyperuniformity in the spatial distribution of points, which is of independent interest in statistical physics [Ghosh2016NumberRI, GhoshLebowitz, PhysRevE]. In fact, for the hyperbolic GAF with $L=1$, it has been shown by Peres and Virag [PeresVirag] that the zeros of $f_{1}$ is a DPP on $\mathbb{D}$ with the Bergman kernel $K(z,w)=\pi^{-1}(1-z\bar{w})^{-2}$ and the background measure is the Lebesgue measure on $\mathbb{D}$. The zeros of GAFs also exhibit repulsiveness, but at short ranges. In particular, the probability of having 2 zeros in a same small disk of radius $\varepsilon$ is of order $\mathcal{O}(\varepsilon^{6})$ (see [hough2009zeros], Chapter 3). In contrast, from (2.2) with $d=2$, that probability for Poisson point processes is of order $\varepsilon^{4}$. For more details about GAFs, we refer the readers to [hough2009zeros, sodin].
@@ 8.2.1 Bargmann transform and Hermite functions
$\mathcal{B}f(z)=2^{1/4}\int_{\mathbb{R}}f(x)\exp\Big(2\pi xz-\pi x^{2}-\frac{\pi}{2}z^{2}\Big.)\mathrm{d}x,\quad z\in\mathbb{C}.$
@@
$\mathcal{F}_{\phi}f(z)=\exp\Big(-\pi iuv-\frac{\pi}{2}|z|^{2}\Big.)\mathcal{B}f(\bar{z}),$
@@ 8.3 The short-time Fourier transforms of white noise
Using Bargmann transform, the STFT of the white noise $\xi$ is then given by the following theorem (see [BARDENET_Spec, GhoshGAF]).
@@
$\mathcal{F}_{\phi}\xi(z)=\sqrt{\pi}\exp\Big(i\pi uv-\frac{\pi}{2}|z|^{2}\Big.)\sum_{k=0}^{\infty}\langle\xi,h_{k}\rangle\sqrt{\frac{\pi^{k}}{k!}}z^{k},$ (8.3)
@@
Since $\langle\xi,h_{k}\rangle$ are i.i.d. real Gaussians, this is also called the symmetric planar Gaussian analytic function. However, the zero set in this case is not stationary, which poses an obstacle in using techniques from spatial statistics. Motivated by early results by Prosen, the authors of [BARDENET_Spec] argue that the zero set of the spectrogram of real Gaussian white noise could be effectively approximated by the zero set of the spectrogram of complex Gaussian white noise. It turns out that the latter is precisely the zero set of the planar GAF, which has been shown to be invariant under isomorphisms of the complex plane.